OpenStudy (anonymous):

What is the slope-intercept form of -3x + 9y = 6?

OpenStudy (anonymous):

9y=3x+6 y=3x+6/9 y=1/3x+2/3 m=1/3 y-intercept=2/3

OpenStudy (anonymous):

THanks